OpenStudy (anonymous):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f(x) = x - 8 / x+7 and g(x) = -7 x - 8 / x -1.

OpenStudy (anonymous):

lots of ugly algebra ending in a raft of cancellation

OpenStudy (anonymous):

\[f(x)=\frac{x-8}{x+7}\] replace \(x\) by \(\frac{-7x-8}{x-1}\) and get \[f(f^{-1}(x))=\frac{\frac{-7x-8}{x-1}-8}{\frac{-7x-8}{x-1}+7}\]

OpenStudy (anonymous):

multiply top and bottom by \(x-1\) to clear the compound fraction then there will be a whole slew of cancellation and you should end up only with \(x\) if the inverse is correct

OpenStudy (anonymous):

I dont understand it still I see what you are trying to do but the print is small on my computer. Can you you do it like one thing at a time ?

OpenStudy (anonymous):

let me make it larger

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

\[\huge f(f^{-1}(x))=\frac{\frac{-7x-8}{x-1}-8}{\frac{-7x-8}{x-1}+7}\]

OpenStudy (anonymous):

So what would be left I think I got it

OpenStudy (anonymous):

if it is all correct, you should be left only with \(\large x\)

OpenStudy (anonymous):

ok I got it thanks

OpenStudy (anonymous):

yw